If a can containing one pound of freon (CCL2F2) were punctured releasing all the contents as a cloud of vapor, what would be the volume of the resulting cloud? V = _________ cubic meters. What would be density of the cloud? Would it float or sink in the air? ________________g/L; _______________.
1
Expert's answer
2012-04-03T11:31:31-0400
1 pound = 0.453 592 37 kilogram and 453.59 g
453.59g v
CCl2F2 --> CCl2F2
121 22.4

v = 453.59 * 22.4/121 = 83.97 L

density is 453.59 / 83.97 = 5.40 g/L

Density of air is 1.293 and density of CCl2F2 is 5.4, so Density of vapor is about 4 times that of air. Vapor will collect in low areas